OpenStudy (anonymous):

Kinematic 1) Terry threw a baseball high above his head and then caught it again before it hits the ground. it total flight time was 5seconds. ignoring air resistance. Calculate the speed at which Terry flung the ball into the air

OpenStudy (anonymous):

Use 1 dimensional equation of motion

OpenStudy (amistre64):

my idea is: Vf = Vi - gt Vi = Vf + gt Vf = 0 when the ball is at its highest point, and it takes half the time to get there. Vi = gt/2 If we were to simple drop the ball, how fast would it be going after 2.5 secs?

OpenStudy (anonymous):

Considering the symmetry of projectile motion you can take the time to reach the maximum height to be half of the total time of flight. So t= 2.5 s. Next a=-g. vf = 0 Now use eqn of motion to find vi.